40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 12 15:14:09 2014, manasa, Update, General, X end AUX laser fiber setup ETMXtable.png
    Reply  Mon Dec 22 17:06:14 2014, manasa, Update, General, X end AUX laser fiber setup 
       Reply  Tue Dec 23 13:18:37 2014, manasa, Update, General, X end AUX laser fiber setup 
          Reply  Wed Jan 7 19:49:09 2015, manasa, Update, General, X end space issues ETMXspace.png
             Reply  Thu Jan 29 10:14:21 2015, manasa, Update, General, X end space issues photo.pngdrawing.png
                Reply  Wed Feb 4 21:09:43 2015, manasa, Update, General, X endtable work scheduled tomorrow 
                   Reply  Thu Feb 5 15:21:25 2015, manasa, Update, General, X endtable work  
                      Reply  Tue Feb 10 16:01:21 2015, manasa, Update, General, AUX X fiber coupled 72% AUXxTelescope.png
Message ID: 10981     Entry time: Thu Feb 5 15:21:25 2015     In reply to: 10978     Reply to this: 10996
Author: manasa 
Type: Update 
Category: General 
Subject: X endtable work  

[EricG, Manasa]

We were at the X end today trying to couple AUX X light into the fiber. 

The proposed plan still did not give a good beampath. The last steering mirror before the fiber coupler was sticking out of the table enclosure. I tried a few other options and the maximum coupling that I could get was ~10%

I am working on plan C now; which would be to use fixed mount mirrors and steer the beam to the space created by Koji near the IR trans path and use a set of lenses instead of a single lens. I will elog more details after some modematching calculations.

We moved one of the clamps for the doubling crystal to make space. Also, the NPRO current was reduced during this work.

I reset things to how they were before I touched the table. I ensured that the green power was still the same (~3mW) after the doubler and that it could lock to the arm in TEM00.

ELOG V3.1.3-